OpenStudy (anonymous):

Consider the second order initial value problem y''+3y'+2y=0, y(0)=1,y'(0)=0 Write this as a system of two first order ordinary differential equations. Make sure you clearly define the dependent variables and state their initial conditions

OpenStudy (dumbcow):

do you have to make it a system? this is homogeneous so just use characteristic equation ---> r^2 +3r+2 = 0

OpenStudy (anonymous):

To solve this, we write it as a quadratic: k^2+3k+2=0 which can be factorised to (k+2)(k+1) giving two roots: k1=-2 and k2=-1. We now get our linearly independent solutions y1=e^(-2x) and y2=e^(-x). Our solution y(x)=Ae^(-2x)+Be^(-x) for constants A and B which is what we need to find now using the given values. y(x)=Ae^(-2x)+Be^(-x) y(0)=1=A+B y'(x)=-2Ae^(-2x)-Be^(-x) y'(0)=0=-2A-B We now have two equations for A and B that we can solve: 1=A+B => B=1-A 0=-2A-B 0=-2A-(1-A) A=-1 and hence B=2. Sub these values back into our equation for y(x) and we get: y(x)=-e^(-2x)+2e^(-x)

OpenStudy (anonymous):

If you let \(u=y^{\prime}\), then it follows that \(u^{\prime} = y^{\prime\prime}\). Thus, \(y^{\prime\prime}+3y^{\prime}+2y = 0\implies u^{\prime}+3u+2y = 0 \implies u^{\prime} = -3u-2y\) The initial condition \(y^{\prime}(0) = 0\) is then equivalent to \(u(0)=0\). Therefore, the original equation can be written as the system of first order ODEs \[\left\{\begin{aligned}y^{\prime} &= u\\ u^{\prime} &= -3u-2y\end{aligned}\right.\] with intial conditions \(y(0)=1\) and \(u(0)=0\). I hope this makes sense! :-)